# Webcam-based Pupil Diameter Prediction Benefits from Upscaling

### v1

@@ 3.2 EyeDentify++
As a result of the examination of GFPGAN, CodeFormer, Real-ESRGAN, HAT, and SRResNet SR models for pupil diameter estimation, we can create five additional datasets containing left and right eye images separately, which we call EyeDentify++. Due to the different SR approximations, the later stages, where we recognize faces, crop eyes, and detect blinks, result in retaining and discarding different amounts of images. More formally, $\lvert\varphi_{blink}\left(\mathbb{D}^{GFPGAN\times 2}_{eyes}\right)\rvert\neq\lvert\varphi_{blink}\left(\mathbb{D}^{HAT\times 2}_{eyes}\right)\rvert$. Figure 3 compares the number of images in the original dataset with those in the SR datasets after blink detection. The results indicate that S enhances the accuracy of blink classification by improving the calculation of the EAR ratio through clearer eye landmark detection on the 2x and 4x up-scaled images and providing higher-quality images for feature extraction in the subsequent blink detection phase [42].
@@ 4.3 Results
Table 1: Quantative Mean Absolute Error (MAE) $\downarrow$ comparison across different pre-trained SR methods and pupil diameter prediction models for both left and right eyes. The lowest errors are highlighted.
@@ 5 Limitations
This study faces several challenges, as shown in Figure 4. Participants were recorded in natural postures with varying distances from the webcam and no strict positioning guidelines, leading to inconsistencies like movement (A), gaze shifts (B), head/body turns (C), and actions like talking or smiling (D). Differences in eye structure, skin tone, and iris color across diverse nationalities and demographics make it difficult to generalize the model. Variations in lighting and screen color changes further affect the perceived eye and pupil colors (E, F, G, H). Additionally, Figure 2 and Figure 4 (A, C, E, G, H) highlight that GAN-based models introduce artifacts like glare, altered eye size, and changes in iris color, complicating training.
@@ 6 Conclusion & Future Work
In this work, we investigated the role of SR techniques in enhancing the accuracy of pupil diameter prediction from webcam-based images, which is crucial for assessing psychological and physiological states. Our experiments, across multiple upscaling methods and neural network architectures, demonstrate that SR can significantly refine the feature details necessary for more precise pupil measurements. Key findings indicate that while the benefits of SR are clear, they are not uniformly distributed across different scales and methods. For instance, although traditional bicubic upscaling often performs well, advanced SR techniques like Real-ESRGAN and SRResNet generally provide superior error rates under specific conditions. In conclusion, while SR presents a promising avenue for enhancing low-quality, webcam-derived images for pupilometry, it requires nuanced application and thorough validation to fully realize its benefits. Future work will aim to extend these initial findings by exploring additional SR methods and integrating more diverse data conditions to ensure the robustness and applicability of pupil diameter estimation techniques in real-world scenarios. This research not only advances our understanding of image upscaling in pupilometry but also sets a strong foundation for future advancements in eye-tracking technologies.
@@ Acknowledgements
Figure 5: Class Activation Map [54] visualizations for the final convolutional layer of ResNet18, ResNet50, and ResNet152 are shown for a test participant viewing the same display color with No-SR, SRx2, and SRx4 eye images. The true and predicted values represent the original and estimated pupil diameters of the left and right eyes in millimeters.

### v2

@@ 3.2 EyeDentify++
As a result of the examination of GFPGAN, CodeFormer, Real-ESRGAN, HAT, and SRResNet SR models for pupil diameter estimation, we can create five additional datasets containing left and right eye images separately, which we call EyeDentify++  [^1] https://vijulshah.github.io/webcam-based-pupil-diameter-estimation/. Due to the different SR approximations, the later stages, where we recognize faces, crop eyes, and detect blinks, result in retaining and discarding different amounts of images. More formally, $\lvert\varphi_{blink}\left(\mathbb{D}^{GFPGAN\times 2}_{eyes}\right)\rvert\neq\lvert\varphi_{blink}\left(\mathbb{D}^{HAT\times 2}_{eyes}\right)\rvert$. Figure 3 compares the number of images in the original dataset with those in the SR datasets after blink detection. The results indicate that S enhances the accuracy of blink classification by improving the calculation of the EAR ratio through clearer eye landmark detection on the 2x and 4x up-scaled images and providing higher-quality images for feature extraction in the subsequent blink detection phase [42].
@@ 4.3 Results
Table 1: Quantitative Mean Absolute Error (MAE) $\downarrow$ comparison across different pre-trained SR methods and pupil diameter prediction models for both left and right eyes. The lowest errors are highlighted.
@@ 5 Limitations
This study faces several challenges, as shown in Figure 4. Participants were recorded in natural postures with varying distances from the webcam and no strict positioning guidelines, leading to inconsistencies like movement (A), gaze shifts (B), head/body turns (C), and actions like talking or smiling (D). Differences in eye structure, skin tone, and iris color across diverse nationalities and demographics make it difficult to generalize the model. Variations in lighting and screen color changes further affect the perceived eye and pupil colors (E, F, G, H). Additionally, Figure 2 and Figure 4 (A, C, E, G, H) highlight that GAN-based models introduce artifacts like glare, altered eye size, and changes in iris color, complicating model training.
@@ 6 Conclusion & Future Work
In this work, we investigated the role of SR techniques in enhancing the accuracy of pupil diameter prediction from webcam-based images, which is crucial for assessing psychological and physiological states. Our experiments, across multiple upscaling methods and neural network architectures, demonstrate that SR can significantly refine the feature details necessary for more precise pupil measurements. Key findings indicate that while the benefits of SR are clear, they are not uniformly distributed across different scales and methods. For instance, although traditional bicubic upscaling often performs well, advanced SR techniques like Real-ESRGAN and SRResNet generally provide superior error rates under specific conditions. In conclusion, while SR presents a promising avenue for enhancing low-quality, webcam-derived images for pupilometry, it requires nuanced application and thorough validation to fully realize its benefits.
@@ Acknowledgements
Figure 5: Class Activation Map [54] visualizations for the final convolutional layer of ResNet18, ResNet50, and ResNet152 are shown for a test participant viewing the same display color with No-SR, SRx2, and SRx4 eye images. The true and predicted values represent the original and estimated pupil diameters.